My teacher is a bad math teacher so I don't get this.
Question: Traveling upstream on the Missouri River, a barge travels 35 mi in 7 h. Downstream, it travels the same distance in half the time. What is the rate of the barge in still water and the rate of the current.
Please use the variable r[rate of barge in still water] and c[rate of current]
Could you please do the 4 steps process. If you do not know what the four step process is..it's 1st: let statement 2nd: equation 3rd: solve 4th: Writing the answer in a complete sentence. Thank you

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

Traveling upstream on the Missouri River, a barge travels 35 mi in 7 h. Downstream, it travels the same distance in half the time. What is the rate of the barge in still water and the rate of the current.
--------------------------------
Upstream DATA:
distance = 35 miles; time = 7 hrs ; rate = d/t = 35/7 = 5 mph
---------------------
Downstream DATA:
distance = 35 miles ; time = (7/2) hrs ; rate = 35/(7/2) = 10 mph
-----
Equations:
b +c = 10
b -c = 5
----
Add and solve for "b":
2b = 15
b = 7 1/2 mph (speed of the barge in still water)
---
Solve for "c":
b + c = 10
7 1/2 + c = 10
c = 2 1/2 mph (speed of the current)
=======================================
Cheers,
Stan H.
==========

Answer by mananth(16949)

Question: Traveling upstream on the Missouri River, a barge travels 35 mi in 7 h. Downstream, it travels the same distance in half the time. What is the rate of the barge in still water and the rate of the current.
Please use the variable r[rate of barge in still water] and c[rate of current]
Could you please do the 4 steps process. If you do not know what the four step process is..it's 1st: let statement 2nd: equation 3rd: solve 4th: Writing the answer in a complete sentence. Thank you
..
r= rate of barge in still water
c= rate of current.
...
downstream rate = r+c
upstream rate = r-c
...
time downstream = 3.5 hours
time upstream = 7 hours
...
d is same both directions = 35 miles
7(r-c)=35
/7
r-c=5......................1
..
3.5(r+c)=35
7/2(r+c)=35
multiply by 2
7(r+c)=70
/7
r+c=10.....................2
add (1) & (2)
2r=15
/2
r=7.5 mph rate of bargein still water
...
r-c=5
7.5-c=5
c=7.5-5
c=2.5 mph rate of current.
..